OpenStudy (anonymous):

Find the difference. m + 2 / (m - 4)^2 - m - 2 / (m - 4)^2 a. 4 / (m - 4)^2 b. 0 c. 2 / (m - 4)^2 d. 2m / (m - 4)^2 please help me.

jimthompson5910 (jim_thompson5910):

The denominators are both (m - 4)^2, so we just subtract the numerators \[\Large \frac{m+2}{(m - 4)^2} + \frac{m-2}{(m - 4)^2}\] \[\Large \frac{(m+2)-(m-2)}{(m - 4)^2}\] \[\Large \frac{m+2-m+2}{(m - 4)^2}\] \[\Large \frac{0m+4}{(m - 4)^2}\] \[\Large \frac{4}{(m - 4)^2}\] so hopefully that's not so bad. Unfortunately, this only works if the denominators were the same. You have to do a bit of more work to get them equal (if they weren't the same).

OpenStudy (anonymous):

thank you , i'll read on this topic a bit more. <3

jimthompson5910 (jim_thompson5910):

good idea, make sure to take a lot of notes as well